OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9 A. (3, 9) B. (–27, –9) C. (–3, –6) D. (–9, –9)

OpenStudy (anonymous):

Are the equations like these: \(3x – 4y = 9\\–3x + 2y = 9\) ? If so, just add the two equations to eliminate for the variable x, then solve for y. After you solve for y, substitute the value that you found for y into any of the equations, then solve for x.

OpenStudy (anonymous):

no its all togather

OpenStudy (anonymous):

like this \(3x – 4y = 9 –3x + 2y = 9\) ?? ಠ_๏ you sure?

OpenStudy (anonymous):

3x – 4y = 9 –3x + 2y = 9 consider 3x – 4y =9 -(1) and 9 –3x + 2y = 9 -(2) (1)+(2) 3x – 4y + ( 9 –3x + 2y) = 9+9 9-2y = 18 2y = 9 y = 9/2 = 4.5 put y= 4.5 into (1) 3x -4(4.5) = 9 3x = 27 x=9